# Robust Universal Inference For Misspecified Models

### v2

@@ Abstract
In statistical inference, it is rarely realistic that the hypothesized statistical model is well-specified, and consequently, it is important to understand the effects of misspecification on inferential procedures. When the hypothesized statistical model is misspecified, the natural target of inference is a projection of the data generating distribution onto the model. We present a general method for constructing valid confidence sets for such projections, under weak regularity conditions, despite possible model misspecification. Our method builds upon the universal inference method and is based on inverting a family of split-sample tests of relative fit. We study settings in which our methods yield either exact or approximate, finite-sample valid confidence sets for various projection distributions. We study rates at which the resulting confidence sets shrink around the target of inference and complement these results with a simulation study.
@@ 2.3 Divergences
and $\mathsf{DP}_{0}=\mathsf{KL}$ is defined by taking the limit of $\beta\to 0$. When $\beta=1$, the density power divergence is the $L_{2}$ distance $\int(p-q)^{2}\,{\mathrm{d}}\lambda$. Finally, the family of Integral Probability Metrics [IPMs, 44], are defined as
@@ 6 Size of Robust Universal Confidence Sets
Theorem 5 also shows how the size of the set depends on the choice of $\widehat{P}_{1}$. When possible we should choose a pilot estimate $\widehat{P}_{1}$ which converges to the target $\widetilde{P}$ at a fast rate to ensure that the term $\eta_{n}$ is sufficiently small. A sensible choice is often a minimum distance estimator [MDE, 45] which is not only a consistent estimator of $\widetilde{P}$ under some regularity conditions but is also robust to some misspecification in its corresponding distance [43, 22].
@@ 7 Illustration: Contaminated Data Generating Process
8 Practical Implementation
  1. 1.
  2. 2.
    For $j=1,\ldots,B$, let
  3. 3.
  4. 4.
    Return $C_{\alpha}=\{\theta:\ \widehat{pv}(\theta)\geq\alpha\}$.
Proposition 8.
@@ E.1 Proof of Theorem 5
Now we focus on bounding $\mathbb{E}_{P^{*}}[\sup_{P\in\mathcal{S}_{\epsilon}}|\widetilde{T}_{P}|]$ (which is greater than $\mathbb{E}_{P^{*}}[\sup_{P\in\mathcal{S}_{\epsilon}}\widetilde{T}_{P}]$). Let $\mathcal{F}_{T,\mathcal{P}}=\{T(\cdot;\widehat{P}_{1},P):P\in\mathcal{P}\}$. The symmetrization lemma [Lemma 2.3.1, 54] states that
@@ I.2 Example 2
Figure 14 compares the various TV ReDI set when the $P^{*}$ is a $32\%$ variance inflated negative binomial—Berry-Esseen ($\widehat{C}_{\mathsf{TV}\textsc{ReDI}}$), Hoeffding bound ($\widehat{C}_{\mathsf{TV}\textsc{HF}}$), empirical Bernstein bound [$\widehat{C}_{\mathsf{TV}\textsc{EBS}}$; 56], and empirical Bentkus bound [$\widehat{C}_{\mathsf{TV}\textsc{EBK}}$; 9, 36]. See Section F for explicit forms of each confidence set. In all cases, we choose the same minimum TV distance estimator $\widehat{\theta}_{1}$. The KL ReDI set dominates all finite sample valid confidence sets considered in this section, despite its validity relying on asymptotics. The finite sample valid ReDI sets are too conservative (and yield a meaningless set $\widehat{C}=\mathcal{P}$) when only a few observations are available ($n\leq 50$). Although our paper does not primarily focus on obtaining the tightest finite-sample valid confidence set, leveraging the variance $\mathbb{V}_{P^{*}}(X)$ can often be beneficial when constructing the confidence set. In this example, $\widehat{C}_{\mathsf{TV}\textsc{EBS}}$ and $\widehat{C}_{\mathsf{TV}\textsc{EBK}}$ outperform $\widehat{C}_{\mathsf{TV}\textsc{HF}}$ since the Bernstein and Bentkus bounds are more sensitive to the variance.

### v3

@@ Abstract
In statistical inference, it is rarely realistic that the hypothesized statistical model is well-specified, and consequently, it is important to understand the effects of misspecification on inferential procedures. When the hypothesized statistical model is misspecified, the natural target of inference is a projection of the data generating distribution onto the model. We present a general method for constructing valid confidence sets for such projections, under weak regularity conditions, despite possible model misspecification. Our method builds upon the universal inference method and is based on inverting a family of split-sample tests of relative fit. We study settings in which our methods yield either exact or approximate, finite-sample valid confidence sets for various projection distributions. We study rates at which the resulting confidence sets shrink around the target of inference and complement these results with a simulation study and a causal discovery using linear causal model on CausalEffectPairs dataset.
@@ 2.3 Divergences
and $\mathsf{DP}_{0}=\mathsf{KL}$ is defined by taking the limit of $\beta\to 0$. When $\beta=1$, the density power divergence is the $L_{2}$ distance $\int(p-q)^{2}\,{\mathrm{d}}\lambda$. Finally, the family of Integral Probability Metrics [IPMs, 45], are defined as
@@ 6 Size of Robust Universal Confidence Sets
Theorem 5 also shows how the size of the set depends on the choice of $\widehat{P}_{1}$. When possible we should choose a pilot estimate $\widehat{P}_{1}$ which converges to the target $\widetilde{P}$ at a fast rate to ensure that the term $\eta_{n}$ is sufficiently small. A sensible choice is often a minimum distance estimator [MDE, 46] which is not only a consistent estimator of $\widetilde{P}$ under some regularity conditions but is also robust to some misspecification in its corresponding distance [43, 22].
7 Practical Implementation
  1. 1.
  2. 2.
    For $j=1,\ldots,B$, let
  3. 3.
  4. 4.
    Return $C_{\alpha}=\{\theta:\ \widehat{pv}(\theta)\geq\alpha\}$.
Proposition 8.
@@ 8 Illustration: Contaminated Data Generating Process
9 Application: Causal Discovery with Homoscedastic Linear Causal Model
In this section, we demonstrate how our proposed method can be applied to construct robust confidence sets for total causal effects using a bivariate Gaussian linear structural equation model (LSEM) with homoscedastic errors [47]. The homoscedastic bivariate Gaussian LSEM is formally defined as follows:
$\mathcal{P}=\left\{\mbox{$\mathsf{N}$}\left(0,\Sigma\right):\Sigma=\sigma^{2}(I_{2}-B)^{-1}(I_{2}-B)^{-\top},\quad B=\begin{pmatrix}0&\beta_{12}\\ \beta_{21}&0\end{pmatrix}\right\},$ (27)
where $\sigma^{2}\in\mathbb{R}_{+}$ is an unknown variance parameter and $\beta_{ij}$ are unknown parameters of direct causal effect from $X_{i}$ to $X_{j}$. Under the model $\mathcal{P}$, the target of inference is the total causal effect $C(i\to j)=(I_{2}-B)^{-1}_{ji}$ for $(i,j)\in\{(1,2),(2,1)\}$, i.e.,
$C(i\to j)=\Sigma_{ij}/\Sigma_{ii}\mathbb{I}(\Sigma\in\mathcal{M}_{r_{i}}),\qquad\mathcal{M}_{r_{i}}:=\left\{\Sigma\succ 0:\Sigma^{2}_{ii}=|\Sigma|\right\}.$
Given iid samples $\bm{X}_{1},\dots,\bm{X}_{n}\sim P^{*}$, our goal is to construct a confidence set for the total causal effect $C(i\to j)$. When the model is well specified ($P^{*}\in\mathcal{P}$), the causal structure is identifiable [47], and one can construct confidence sets by inverting either (A) likelihood ratio tests that specify $C(i\to j)$, with the alternative relaxed to the positive definite cone instead of the union $\mathcal{M}{r_{1}}\cup\mathcal{M}{r_{2}}$, or (B) the split likelihood ratio test [53]. Strieder et al. [53] showed that the LRT can be sensitive to model misspecification despite having narrower confidence set than the split LRT confidence set. Recall that split LRT is valid only when model is well specified (or convex) and its implicit target of inference— the KL projection—can be sensitive to the model misspecification. Therefore we compare the LRT set, the split LRT set, the $\mathsf{KL}$ReDI, and the $\mathsf{DP}$ReDI sets including crossfit variants on datasets exhibiting varying degrees and types of model misspecification. See Supplementary Material L for details on constructing ReDI sets under homoscedastic bivariate Gaussian LSEM, and for additional simulation results.
We used the CausalEffectPairs [44] dataset, which offers a collection of cause-effect pairs from diverse domains, with ground-truth causal directions established through domain expertise. For our application, we consider three representative pairs: pair 76 (the average annual growth of population $X_{1}$ and the food consumption $X_{2}$), pair 18 (age of child $X_{1}$ and chemical GAG concentration in urine $X_{2}$), and pair 8 (abalone age $X_{1}$ and height $X_{2}$). These examples were selected to illustrate varying degrees and types of model misspecification. Pair 76 is considered nearly well-specified and was also analyzed in Strieder et al. [53]. Pair 18 exhibits nonlinearity and possible deviations from homoscedasticity, while pair 8 involves both nonlinearity and heteroscedasticity, as well as notable outliers. See Figure 6 for scatterplots of each pairs after centering.
Figure 6: Scatterplots of selected CausalEffectPairs after centering.
Figure 7: 95% confidence intervals for the total causal effect from $X_{1}$ to $X_{2}$ (Top) and from $X_{2}$ to $X_{1}$ (Bottom) for selected CausalEffectPairs.
Figure 7 shows the 95% confidence sets for the total causal effects $C(1\to 2)$ and $C(2\to 1)$, constructed using two LRT variants [LRT1 and LRT2, 53], split LRT, KLReDI, DPReDI with $\beta=0.1,1$. For split LRT and ReDI sets, we provide both vanilla variant—based on equally split data—and the crossfit variant, which averages the split test statistics and its corrupted variances by swapping the role of split samples $\mathcal{D}_{0}$ and $\mathcal{D}_{1}$. This crossfitting approach can help reduce split-sample variability and improve efficiency. See Supplementary Material G for details. All methods correctly identified the ground-truth causal direction across all three pairs. LRT2 fails to provide confidence set in all cases, despite its strong simulation performance, whereas LRT1 yielded the narrowest confidence intervals.
In the nearly well-specified example (e.g., pair 76), the split LRT and the KLReDI produced comparable confidence sets. Among all methods, $\mathsf{DP}\textsc{ReDI}_{1}$ yielded the widest interval, consistent with the lower efficiency of minimum density power divergence methods compared to MLE [7]. Interestingly, $\mathsf{DP}\textsc{ReDI}_{0.1}$ resulted in a narrower confidence interval than both KLReDI and split LRT. This suggests that $\mathsf{DP}\textsc{ReDI}$ may have effectively downweighted influential observations (e.g., $(X_{1},X_{2})=(-1.6,9.5)$), achieving robustness with only a modest loss in efficiency.
For pair 18, where the model is moderately misspecified, the KLReDI produced wider confidence sets than the split LRT, appropriately reflecting the uncertainty due to misspecification. In this case, $\mathsf{DP}\textsc{ReDI}$ (for both $\beta=0.1$ and $\beta=1$) yielded even wider sets than KLReDI set. As $\beta$ increases, the $\mathsf{DP}\textsc{ReDI}$ confidence set shifts slightly toward larger values of $C(1\to 2)$, though the shift remains minor. This shift reflects the effect of DP divergence downweighting influential observations near $X_{1}=-5$,
For pair 8, where the data exhibit nonlinearity, heteroscedasticity, and contain outliers, KLReDI produces significantly wider confidence sets than split LRT, reflecting the impact of model misspecification. In contrast, the $\mathsf{DP}\textsc{ReDI}$ set is comparable in width to, or even narrower than, split LRT. This highlights the sensitivity of KL projection to small but adversarial perturbations (as discussed in Example 3 and Section 8), and underscores the benefit of targeting a more stable projection for inference. The influence of the DP parameter $\beta$ is especially clear in this example: as $\beta$ increases, the confidence set for $C(2\to 1)$ shifts toward larger values. This is due to the increased robustness of the DP projection, which more effectively downweights outliers (e.g., $(X_{1},X_{2})=(-1.9,1.0)$). Finally, note that the LRT1 and cross-fit $\mathsf{DP}\textsc{ReDI}_{1}$ sets do not overlap, indicating that inference based on an overconfident LRT—without accounting for model misspecification—can substantially underestimate the total causal effect $C(2\to 1)$.
@@ E.1 Proof of Theorem 5
Now we focus on bounding $\mathbb{E}_{P^{*}}[\sup_{P\in\mathcal{S}_{\epsilon}}|\widetilde{T}_{P}|]$ (which is greater than $\mathbb{E}_{P^{*}}[\sup_{P\in\mathcal{S}_{\epsilon}}\widetilde{T}_{P}]$). Let $\mathcal{F}_{T,\mathcal{P}}=\{T(\cdot;\widehat{P}_{1},P):P\in\mathcal{P}\}$. The symmetrization lemma [Lemma 2.3.1, 57] states that
@@ I.2 Example 2
Figure 16 compares the various TV ReDI set when the $P^{*}$ is a $32\%$ variance inflated negative binomial—Berry-Esseen ($\widehat{C}_{\mathsf{TV}\textsc{ReDI}}$), Hoeffding bound ($\widehat{C}_{\mathsf{TV}\textsc{HF}}$), empirical Bernstein bound [$\widehat{C}_{\mathsf{TV}\textsc{EBS}}$; 59], and empirical Bentkus bound [$\widehat{C}_{\mathsf{TV}\textsc{EBK}}$; 9, 36]. See Section F for explicit forms of each confidence set. In all cases, we choose the same minimum TV distance estimator $\widehat{\theta}_{1}$. The KL ReDI set dominates all finite sample valid confidence sets considered in this section, despite its validity relying on asymptotics. The finite sample valid ReDI sets are too conservative (and yield a meaningless set $\widehat{C}=\mathcal{P}$) when only a few observations are available ($n\leq 50$). Although our paper does not primarily focus on obtaining the tightest finite-sample valid confidence set, leveraging the variance $\mathbb{V}_{P^{*}}(X)$ can often be beneficial when constructing the confidence set. In this example, $\widehat{C}_{\mathsf{TV}\textsc{EBS}}$ and $\widehat{C}_{\mathsf{TV}\textsc{EBK}}$ outperform $\widehat{C}_{\mathsf{TV}\textsc{HF}}$ since the Bernstein and Bentkus bounds are more sensitive to the variance.
@@
Appendix L Additional Results and Technical Details of Section 9
L.1 ReDI sets for LSEM
KL ReDI Recall that KLReDI utilizes log likelihood ratios as in sLRT. We follow the same strategy of profile sLRT with profile likelihood ratio and use the (corrupted) estimated variance of profile log likelihood ratios to set the critical value. Formally, suppose we observe iid $n$ observations $\mathcal{D}:=\{\bm{X}_{1},\ldots\bm{X}_{n}\}\sim{P^{*}}^{\otimes n}$. Randomly split $\mathcal{D}$ into two equally sized data $\mathcal{D}_{0}$ and $\mathcal{D}_{1}$. Denote their sample sizes $n_{0}$ and $n_{1}$, respectively. By LSEM construction 27, we have the Gaussian log likelihood as follows:
$\ell(\Sigma;\bm{X}_{i})=-\log(2\pi)-\frac{1}{2}\log|\Sigma|-\frac{1}{2}\bm{X}_{i}^{\top}\Sigma^{-1}\bm{X}_{i}.$
Since the functional of interest is $\psi:=C(i\to j)$ and thus $\sigma^{2}$ is a nuisance parameter, we profile out the $\sigma$ for a given $\psi$, i.e.,
$\ell_{\dagger}(\psi;\bm{X}_{i}):=\ell(\widehat{\sigma}^{2}_{0,\psi}\Lambda;X_{i})=-\log(2\pi)-\frac{1}{2}\log\widehat{\sigma}^{2}_{0,\psi}|\Lambda|-\frac{1}{2\widehat{\sigma}^{2}_{0,\psi}}\bm{X}_{i}^{\top}\Lambda^{-1}\bm{X}_{i},\qquad\widehat{\sigma}^{2}_{0,\psi}=\operatorname*{arg\,max}_{C(i\to j)=\psi}\ell(\Sigma;\mathcal{D}_{0}),$
where $\Lambda\equiv\Lambda(\psi)$ is a structured correlation matrix depends on $\psi$. Conveniently, $\widehat{\sigma}^{2}_{0,\psi}$ has an analytic solution for Gaussian likelihood:
$\widehat{\sigma}^{2}_{0,\psi}=\operatorname{tr}\left({\Lambda(\psi)^{-1}S_{0}}\right)/2,\quad S_{0}=\frac{1}{n_{0}}\sum_{\bm{X}_{i}\in\mathcal{D}_{0}}\bm{X}_{i}\bm{X}_{i}^{\top}.$
Let us detail $\Lambda(\psi)$. When $\psi\neq 0$, $\Sigma\in\mathcal{M}_{r1}$ and thus
$\Lambda(\psi)=\begin{pmatrix}1&\psi\\ \psi&1+\psi^{2}\end{pmatrix}.$ (37)
When $\psi=0$, $\Sigma$ can be any matrix in $\mathcal{M}_{r2}$ and thus
$\Lambda(0)=B(\widetilde{\xi})=\begin{pmatrix}1+\widetilde{\xi}^{2}&\widetilde{\xi}\\ \widetilde{\xi}&1\end{pmatrix}\text{ s.t. }\widetilde{\xi}=\operatorname*{arg\,min}_{\xi}\sum_{\bm{X}_{i}\in\mathcal{D}_{0}}\bm{X}_{i}^{\top}B(\xi)^{-1}\bm{X}_{i}=(S_{0})_{(i,j)}/(S_{0})_{(j,j)},$ (38)
since $|B(\xi)|=1$ for any $\xi$. Then the test statistic of profile KL ReDI is as follows:
$T_{i}(P_{\psi},\widehat{P}_{1})=\ell_{\dagger}(\psi;\bm{X}_{i})-\ell(\widehat{\Sigma}_{1};\bm{X}_{i})=-\frac{1}{2}\left(\log\widehat{\sigma}^{2}_{0,\psi}|\Lambda(\psi)|-\log|\widehat{\Sigma}_{1}|+\bm{X}_{i}^{\top}(\Lambda^{-1}(\psi)/\widehat{\sigma}^{2}_{0,\psi}-\widehat{\Sigma}_{1}^{-1})\bm{X}_{i}\right),$
where $\widehat{\Sigma}_{1}$ is the pilot estimator of $\Sigma$ with $\mathcal{D}_{1}$. A sensible choice of pilot estimator is MLE constrained in $\mathcal{M}=\mathcal{M}_{r1}\cup\mathcal{M}_{r2}$:
$\widehat{\Sigma}_{1}=\operatorname*{arg\,max}_{\Sigma\in\mathcal{M}}\ell(\Sigma;\mathcal{D}_{1})=\operatorname*{arg\,min}_{\Sigma\in\mathcal{M}}\left(\log|\Sigma|+\operatorname{tr}\left({\Sigma^{-1}S_{1}}\right)\right),\quad S_{1}=\frac{1}{n_{1}}\sum_{\bm{X}_{i}\in\mathcal{D}_{1}}\bm{X}_{i}\bm{X}_{i}^{\top}.$
Finally we define the corresponding profile KL ReDI set for $C(i\to j)$ as
$\widehat{C}_{\mathsf{KL}\textsc{ReDI},n}=\left\{\psi\in\mathbb{R}:\overline{T}_{n_{0},\delta}(P_{\psi},\widehat{P}_{1})\leq\frac{z_{\alpha}\widehat{s}_{P,\delta}}{\sqrt{n_{0}}}\right\}.$ (39)
DP ReDI Given that DP ReDI statistic is the plug-in estimator of the difference in DP divergences, we adopt the similar profiling as we did for profile KL ReDI set 39. Formally, the empirical DP divergence is proportional to the following:
$\mathsf{DP}_{\beta}(\mathbb{P}_{n_{0}}\|P)$ $\propto\int p^{1+\beta}\,{\mathrm{d}}\lambda-\left(1+\frac{1}{\beta}\right)\frac{1}{n_{0}}\sum_{\bm{X}_{i}\in\mathcal{D}_{0}}p^{\beta}(\bm{X}_{i})$
$=\frac{(2\pi)^{-\beta}}{|\Sigma|^{\beta/2}}\left[\frac{1}{1+\beta}-\left(1+\frac{1}{\beta}\right)\frac{1}{n_{0}}\sum_{\bm{X}_{i}\in\mathcal{D}_{0}}\exp\left(-\frac{\beta}{2}\bm{X}_{i}^{\top}\Sigma^{-1}\bm{X}_{i}\right)\right]:=\frac{1}{n_{0}}\sum_{\bm{X}_{i}\in\mathcal{D}_{0}}e\mathsf{DP}_{\beta}(\Sigma;\bm{X}_{i}),$
where the last equality is from the fact that $P=\mbox{$\mathsf{N}$}(0,\Sigma)$.
Given that $\Sigma=\sigma^{2}\Lambda(\psi)$ s.t. $|\Lambda|=1$,
$e\mathsf{DP}_{\beta}(\sigma^{2}\Lambda(\psi);\mathcal{D}_{0})=\frac{1}{(2\pi)^{\beta}\sigma^{2\beta}}\left[\frac{1}{1+\beta}-\left(1+\frac{1}{\beta}\right)\frac{1}{n_{0}}\sum_{\bm{X}_{i}\in\mathcal{D}_{0}}\exp\left(-\frac{\beta}{2\sigma^{2}}\bm{X}_{i}^{\top}\Lambda^{-1}\bm{X}_{i}\right)\right]$
Let us profile out the $\sigma$ for a given $\psi$, i.e.,
$e\mathsf{DP}_{\beta,\dagger}(\psi;\bm{X}_{i})=e\mathsf{DP}_{\beta}\left(\widehat{\sigma}^{2}_{0,\psi}\Lambda(\psi);\bm{X}_{i}\right),\qquad\widehat{\sigma}^{2}_{0,\psi}=\operatorname*{arg\,min}_{\sigma:C(i\to j)=\psi}e\mathsf{DP}_{\beta}(\sigma^{2}\Lambda(\psi);\mathcal{D}_{0}).$
When $\psi\neq 0$, $\Lambda$ is the same as (37) and thus solving above minimization suffices. When $\psi=0$, $\Sigma$ can be any matrix in $\mathcal{M}_{r2}$ and thus has the same form as (38) except that
$\widetilde{\xi}=\operatorname*{arg\,min}_{\xi}e\mathsf{DP}_{\beta}(\widehat{\sigma}^{2}_{0,0}B(\xi);\mathcal{D}_{0}).$
Unlike $\mathsf{KL}$ReDI, $\widehat{\sigma}^{2}_{0,\psi}$ is analytically intractable and thus we iteratively update $\widehat{\sigma}^{2}_{0,0}$ and $\widetilde{\xi}$ using numerical optimization (i.e., fixed-point iteration).
Then we define the profile $\mathsf{DP}$ ReDI test statistic as follows:
$T_{i}(P_{\psi},\widehat{P}_{1})=e\mathsf{DP}_{\beta,\dagger}(\psi;\bm{X}_{i})-e\mathsf{DP}_{\beta}(\widehat{\Sigma}_{1};\bm{X}_{i}),$
and construct profile DP ReDI set for $C(i\to j)$ analogous to (39).
For the pilot estimator, it is sensible to choose minimum DP estimator constrained in $\mathcal{M}$:
$\operatorname*{arg\,min}_{\Sigma\in\mathcal{M}}e\mathsf{DP}_{\beta}(\Sigma;\mathcal{D}_{1}).$
L.2 Simulation
We generate $n\in\{100,\ldots,1000\}$ i.i.d samples following the model (27) with $\sigma^{2}=1,\beta_{21}=0$ and thus $C(1\to 2)=\beta_{12}\in\{0,0.1,0.2,0.5\}$. We then construct 95% confidence intervals for $C(1\to 2)$ using 6 different methods: LRT1, LRT2, sLRT, KLREDI, and DPREDI (with $\beta=0.1$ and $1$). We also construct both crossfit and vanila sLRT, KLREDI, and DPREDI set, respectively. For each methods, we assess the empirical coverage by the average number of times that $C(1\to 2)$ is included in the confidence interval and average maximum width out of 500 replications.
Figure 21 shows the empirical coverage of each method. All methods yield valid coverage except the LRT1 slightly under covering in $n=500,C(1\to 2)=0.1$. Compared to LRT1, LRT2, both universal inference and robust universal inference methods shows conservative coverage.
Figure 21: Empirical coverage by varying $C(1\to 2)$ and sample sizes in well-specified setting
Figure 22 shows the average maximum width of the confidence set. Here we denote maximum width since the resulting confidence set may be the union of 00{0} and some interval not including 00. The width of confidence set the largest for DPReDI with $\beta=1$, followed by DPReDI with $\beta=0.1$, KLReDI, sLRT, LRT1 and LRT2. This ordering underscores the conservatism of (robust) universal inference method as well as of choosing robust target of inference—The higher the $\beta$ is, the more robust the DP projection becomes—when the model is well specified. In all (robust) universal inference methods, crossfit reduced the size of the confidence set without sacrificing the validity. Particularly, the width of the crossfit KLReDI is comparable to vanilla sLRT. This result precisely matches with the power comparison under Gaussian location model in the Supplementary Material S8.1.
Figure 22: Average maximum width by varying $C(1\to 2)$ and sample sizes in well-specified setting